OpenStudy (anonymous):

what is a common multiple of 12 and 15

OpenStudy (anonymous):

60 find the lcm.

OpenStudy (ikram002p):

common multiple is the number that can devide both 12 & 15 so there is infinitly many multiple , any way u shud found the least comon multiple (LCM) that devide both like this 12=2*2*3 15=3*5 as u can see greatest comon devisor(GCD) is 3 LCM=12*15/GCD